# Joint Distillation for Fast Likelihood Evaluation and Sampling in Flow-based Models

### v1

@@ Abstract
Log-likelihood evaluation enables important capabilities in generative models, including model comparison, certain fine-tuning objectives, and many downstream applications. Yet paradoxically, some of today’s best generative models – diffusion and flow-based models – still require hundreds to thousands of neural function evaluations (NFEs) to compute a single likelihood. While recent distillation methods have successfully accelerated sampling to just a few steps, they achieve this at the cost of likelihood tractability: existing approaches either abandon likelihood computation entirely or still require expensive integration over full trajectories. We present fast flow joint distillation (F2D2), a framework that simultaneously reduces the number of NFEs required for both sampling and likelihood evaluation by two orders of magnitude. Our key insight is that in continuous normalizing flows, the coupled ODEs for sampling and likelihood are computed from a shared underlying velocity field, allowing us to jointly distill both the sampling trajectory and cumulative divergence using a single model. F2D2 is modular, compatible with existing flow-based few-step sampling models, and requires only an additional divergence prediction head. Experiments demonstrate F2D2’s capability of achieving accurate log-likelihood with few-step evaluations while maintaining high sample quality, solving a long-standing computational bottleneck in flow-based generative models. As an application of our approach, we propose a lightweight self-guidance method that enables a 2-step MeanFlow model to outperform a 1024 step teacher model with only a single additional backward NFE.
@@ 1 Introduction
Log-likelihood evaluation and likelihood-based inference have long been fundamental to statistical modeling and machine learning, serving as the backbone for parameter estimation (fisher1922foundations), model selection (akaike1974aic), and hypothesis testing (neyman1933tests). In the era of generative AI, the ability to efficiently evaluate log-likelihood (log-density) has become even more critical, as it directly enables key post-training techniques including reinforcement learning and preference optimization, where likelihoods are important for methods like PPO, DPO and GRPO (schulman2017proximal; ouyang2022training; rafailov2023direct; shao2024deepseekmath). Beyond these applications, optimizing log-likelihood also encourages generative models to capture all modes of the data distribution, avoiding mode collapse that plagues adversarial approaches (razavi2019generating).
@@
While likelihood evaluation is useful for modern generative modeling, the most successful generative models for images and video (rombach2022high; flux2025kontext; openai2024sora; polyak2024moviegen; google2025veo3), namely diffusion and flow matching models, suffer from a critical weakness: computing likelihood requires prohibitively expensive iterative neural function evaluations (NFEs). In particular, discrete-time diffusion models like DDPM (ho2020denoising; nichol2021improved) require summing up variational bounds across all timesteps, which needs hundreds to thousands of forward passes to compute a single likelihood. Similarly, continuous-time formulations like score SDE (song2020score) and flow matching (lipman2022flow; albergo2022building; liu2022flow; albergo2023stochasticinterpolantsunifyingframework) also must integrate the divergence along the learned (probability) flow ODE, which typically requires numerical integration with 100-1000 NFEs for accurate likelihood evaluation. While advanced solvers can significantly reduce NFEs (karras2022elucidating), they fundamentally cannot escape the integration requirement and produce vastly inaccurate results when restricted to very few steps ($\leq$ 10 NFEs). This computational burden makes many likelihood-based finetuning objectives, model comparison, and downstream applications prohibitively expensive for modern diffusion/flow matching models.
@@
Interestingly, diffusion and flow matching models faced the same NFE bottleneck for sampling when they were first introduced, where they initially required 1000+ steps to generate a single image. Research addressing this issue has been remarkably successful, with methods that learn to skip multiple steps, either through distillation or self-consistency objectives, emerging as particularly powerful solutions (salimans2022progressive; song2023consistency; kim2023consistency; frans2024one; geng2025mean; boffi2025flowmapmatchingstochastic; boffi2025buildconsistencymodellearning). However, despite achieving few-step sampling, most of these methods completely lose the ability to compute likelihoods, and the few methods that preserve likelihood computation (e.g. kim2023consistency) still require integrating over the entire trajectory with hundreds of NFEs, making fast likelihood evaluation impossible. Thus, while practitioners have solutions to fast sampling for diffusion and flow matching models, fast log-likelihood evaluation remains an unsolved problem. Here, we show that it is possible to achieve both at the same time.
@@
Contributions. Our key insight is that, in flow matching and continuous normalizing flows (CNFs) in general (chen2018neural), computing exact likelihoods requires solving coupled ODEs: the sampling trajectory $\frac{\mathrm{d}}{\mathrm{d}t}x_{t}=v_{\theta}(x_{t},t)$ and the log-density evolution $\frac{\mathrm{d}}{\mathrm{d}t}\log p_{t}(x_{t})=-\operatorname{div}(v_{\theta}(x_{t},t))$, both depending on the same learned velocity $v_{\theta}$. Since the divergence term can be viewed as another output derived from the same velocity model, we can learn to distill both the flow trajectory and its corresponding divergence computation simultaneously within a single model. By jointly optimizing for both accurate few-step sampling and log-likelihood evaluation, we can potentially achieve a model that succeeds at both tasks.
@@
Based on these insights, we propose fast flow joint distillation (F2D2), a simple and modular framework for jointly learning fast sampling and fast log-likelihood evaluation in flow-based models. Our key idea is to leverage the flow map framework (boffi2025buildconsistencymodellearning) to train a single model to predict both the sampling trajectory and cumulative divergence in parallel using a joint self-distillation objective, ensuring both outputs learn to skip the numerous steps in training. This makes F2D2 plug-and-play with any CNF-based few-step sampling method like shortcut models and MeanFlow, and requires only a new divergence prediction head alongside the existing velocity prediction. To our knowledge, F2D2 is the first method to enable accurate few-step log-likelihood evaluation in diffusion/CNF-based generative models, solving a long-standing limitation of these frameworks.
@@
We demonstrate that our method produces both calibrated likelihoods and high quality samples with few-step NFEs on image datasets CIFAR-10 (cifar10) and ImageNet $64\times 64$ (imagenet). We show that our F2D2 are compatible with and can be directly apply to pre-trained shortcut models, MeanFlow and a new distillation method we propose in this paper.
@@
As an application of our method, we introduce maximum likelihood self-guidance, a lightweight test-time intervention which uses rapid likelihood evaluation to optimize over generated samples, requiring only an additional forward and backward pass through the model. Remarkably, we show that F2D2 with maximum likelihood self-guidance instantiated with 2-step MeanFlow achieves lower FID than a 1024-step flow matching model of the same size on CIFAR-10. This proof of concept demonstrates the expanded algorithmic sandbox enabled by rapid likelihood evaluation.
@@ 2 Background
Flow Matching. Flow matching (lipman2022flow; albergo2022building; liu2022flow; albergo2023stochasticinterpolantsunifyingframework) is a scalable training method for generative modeling that learns a time-dependent velocity field $v_{\theta}:\mathbb{R}^{d}\times[0,1]\rightarrow\mathbb{R}^{d}$ to transport samples from a simple noise distribution $p_{0}$ to the data distribution $p_{1}$. Along a straight-line path $x_{t}=(1-t)x_{0}+tx_{1}$ that linearly interpolates between a noise sample $x_{0}\sim p_{0}$ and a data sample $x_{1}\sim p_{1}$, it models the evolution dynamic with the ordinary differential equation ODE:
@@
This objective encompasses diffusion models as a special case with different interpolation schemes (gao2025diffusionmeetsflow). For sampling, it solves the ODE from $t=0$ to $t=1$ with numerical solvers like Euler or $\operatorname{dopri5}$, and typically requires 100-1000 NFEs for high quality samples.
Continuous Normalizing Flows and Likelihood Computation. A flow-matching model is a special case of a continuous normalizing flow (CNF) (chen2018neural), which transports data from an initial distribution $x_{0}\sim p_{0}$ to an estimated distribution by integrating an ODE. In the case of flow-based models, this is precisely Eq.˜2.1. An advantage of the CNF formalism is the ability to explicitly compute likelihoods via the coupled system of ODEs:
@@
Likelihood evaluation is typically expensive, requiring both careful, finely-discretized integration of an ODE across time steps, and a computation of the divergence term whose exact computation (or the variance of its randomized estimator (grathwohl2018ffjord)) scales at least linearly in ambient dimension $d$. Thus, likelihood evaluation is far more computationally burdensome than sampling.
Few-Step Flow-based Models. To address the computational expense of multiple ODE integrations in sampling, recent few-step flow-based models (kim2023consistency; frans2024one; geng2025mean; boffi2025buildconsistencymodellearning; boffi2025flowmapmatchingstochastic) learn to directly predict the outcome of integrating the ODE in Eq.˜2.1 using only a small number of function evaluations (NFEs). These methods can be viewed as sharing a common strategy of learning to predict the flow map of the underlying ODE.
@@ 3 Method
Our key insight is that we can apply few-step flow-based models to the ODE in Eq.˜2.3 which jointly parametrizes sampling and likelihood evaluation. Following boffi2025buildconsistencymodellearning; frans2024one; geng2025mean, we adopt a linear parametrization of the flow map:
@@
Proposition 3.1 (Flow Map Conditions (boffi2025buildconsistencymodellearning)).
@@
Theoretical Justification. To justify the parameterization Eq.˜3.3, we recall a property denoted the tangent condition by boffi2025buildconsistencymodellearning, leveraged by kim2023consistency; geng2025mean to recover the instantaneous velocity and divergence in the $s\to t$ limit:
@@ 3.2 Instantiating F2D2 with Shortcut and MeanFlow
Though our method is, in principle, compatible with any flow map-based method, we instantiate our formulation for Shortcut Models (frans2024one), based on the semigroup property, and for MeanFlow (geng2025mean), based on the Eulerian equation.
@@ 3.2.1 Joint Shortcut: Shortcut-F2D2
Shortcut models (frans2024one) enforce the semigroup property (Proposition˜3.1 (c)) by applying it to the midpoint between timesteps $t$ and $s$. This amounts to the shortcut self-consistency loss:
@@ 3.2.2 Joint MeanFlow: MeanFlow-F2D2
Alternatively, MeanFlow (geng2025mean) learns the time-averaged velocity $u_{\theta}(x,t,s)$ by enforcing the so-called MeanFlow identity $u_{\theta}(x_{t},s,t)=v(x_{t},t)+(s-t)\,\frac{d}{dt}u_{\theta}(x_{t},s,t)$, which we show in Corollary B.2 in Appendix B solves the Eulerian equation (Proposition˜3.1 (b)). To our knowledge, this is the first proof of this fact.
@@ 3.3 Practical Design Choices
Hutchinson Trace Estimator. Computing divergence terms $\operatorname{div}(v(x_{t},t))$ requires $O(d)$ backward passes for exact computation, making training prohibitively expensive for high-dimensional data. Following standard practice (grathwohl2018ffjord; lipman2022flow; song2020score), we employ the Hutchinson trace estimator $\operatorname{div}(v)\approx\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}[\epsilon^{\top}\nabla_{x}v\cdot\epsilon]$ which provides unbiased estimates with only $O(1)$ computational cost per training step.
Staged Training with Warm Start. Since divergence supervision depends on having accurate velocity predictions, we adopt a staged training approach. In practice, we pre-train the sampling velocity component $u_{\theta}$ alone using existing flow map distillation techniques, which provides a good initialization for joint training later. Optionally, we can also pre-train a teacher flow matching model $v_{\phi}$ that serves as a reliable source of divergence supervision, replacing the potentially noisy $\operatorname{div}(u_{\theta}(x_{t},t,t))$ with the more accurate $\operatorname{div}(v_{\phi}(x_{t},t))$ during joint distillation.
@@ 3.4 Application: Maximum Likelihood Self-Guidance with F2D2
Now that we have access to log-likelihood computation with few NFEs, we can explore various applications. One particularly interesting one is using the one-step divergence prediction (combined with the source distribution’s log-likelihood) as a pseudo-likelihood objective for inference-time optimization. Specifically, we can optimize the initial noise $x_{0}$ to improve sample quality before running the sampling procedure. This approach resembles reward-based initial noise optimization for one-step generation models (eyring2024reno), except we do not require external reward models. Instead, we obtain the guidance signal from the model’s own likelihood prediction head – effectively performing self-guidance at inference time to improve sample quality. This maximum likelihood self-guidance sampling algorithm is described in Algorithm 2.

### v2

@@ Abstract
Log-likelihood evaluation enables important capabilities in generative models, including model comparison, certain fine-tuning objectives, and many downstream applications. Yet paradoxically, some of today’s best generative models – diffusion and flow-based models – still require hundreds to thousands of neural function evaluations (NFEs) to compute a single likelihood. While recent distillation methods have successfully accelerated sampling to just a few steps, they achieve this at the cost of likelihood tractability: existing approaches either abandon likelihood computation entirely or still require expensive integration over full trajectories. We present fast flow joint distillation (F2D2), a framework that simultaneously reduces the number of NFEs required for both sampling and likelihood evaluation by two orders of magnitude. Our key insight is that in continuous normalizing flows, the coupled ODEs for sampling and likelihood are computed from a shared underlying velocity field, allowing us to jointly distill both the sampling trajectory and cumulative divergence using a single model. F2D2 is modular, compatible with existing flow-based few-step sampling models, and requires only an additional divergence prediction head. Experiments demonstrate F2D2’s capability of achieving accurate log-likelihood with few-step evaluations while maintaining high sample quality, solving a long-standing computational bottleneck in flow-based generative models. As an application of our approach, we propose a lightweight self-guidance method that enables a 2-step MeanFlow to outperform a 1024 step flow matching model with only a single additional backward NFE.
@@ 1 Introduction
Log-likelihood evaluation and likelihood-based inference have long been fundamental to statistical modeling and machine learning, serving as the backbone for parameter estimation (Fisher, 1922), model selection (Akaike, 1974), and hypothesis testing (Neyman & Pearson, 1933). In the era of generative AI, the ability to efficiently evaluate log-likelihood (log-density) has become even more critical, as it directly enables key post-training techniques including reinforcement learning and preference optimization, where likelihoods are important for methods like PPO, DPO and GRPO (Schulman et al., 2017; Ouyang et al., 2022; Rafailov et al., 2023; Shao et al., 2024). Beyond these applications, optimizing log-likelihood also encourages generative models to capture all modes of the data distribution, avoiding mode collapse that plagues adversarial approaches (Razavi et al., 2019).
@@
While likelihood evaluation is useful for modern generative modeling, the most successful generative models for images and video (Rombach et al., 2022; Black Forest Labs, 2025; OpenAI, 2024; Polyak et al., 2024; Google DeepMind, 2025), namely diffusion and flow matching models, suffer from a critical weakness: computing likelihood requires prohibitively expensive iterative neural function evaluations (NFEs). In particular, discrete-time diffusion models like DDPM (Ho et al., 2020; Nichol & Dhariwal, 2021) require summing up variational bounds across all timesteps, which needs hundreds to thousands of forward passes to compute a single likelihood. Similarly, continuous-time formulations like score SDE (Song et al., 2020) and flow matching (Lipman et al., 2022; Albergo & Vanden-Eijnden, 2022; Liu et al., 2022; Albergo et al., 2023) also must integrate the divergence along the learned (probability) flow ODE, which typically requires numerical integration with 100-1000 NFEs for accurate likelihood evaluation. While advanced solvers can significantly reduce NFEs (Karras et al., 2022), they fundamentally cannot escape the integration requirement and produce vastly inaccurate results when restricted to very few steps ($\leq$ 10 NFEs). This computational burden makes many likelihood-based finetuning objectives, model comparison, and downstream applications prohibitively expensive for modern diffusion/flow matching models.
@@
Interestingly, diffusion and flow matching models faced the same NFE bottleneck for sampling when they were first introduced, where they initially required 1000+ steps to generate a single image. Research addressing this issue has been remarkably successful, with methods that learn to skip multiple steps, either through distillation or self-consistency objectives, emerging as particularly powerful solutions (Salimans & Ho, 2022; Song et al., 2023; Kim et al., 2023; Frans et al., 2024; Geng et al., 2025; Boffi et al., 2025b; a). However, despite achieving few-step sampling, most of these methods completely lose the ability to compute likelihoods, and the few methods that preserve likelihood computation (e.g. Kim et al. (2023)) still require integrating over the entire trajectory with hundreds of NFEs, making fast likelihood evaluation impossible. Thus, while practitioners have solutions to fast sampling for diffusion and flow matching models, fast log-likelihood evaluation remains an unsolved problem. Here, we show that it is possible to achieve both at the same time.
@@
Contributions. Our key insight is that, in flow matching and continuous normalizing flows (CNFs) in general (Chen et al., 2018), computing exact likelihoods requires solving coupled ODEs: the sampling trajectory $\frac{\mathrm{d}}{\mathrm{d}t}x_{t}=v_{\theta}(x_{t},t)$ and the log-density evolution $\frac{\mathrm{d}}{\mathrm{d}t}\log p_{t}(x_{t})=-\operatorname{div}(v_{\theta}(x_{t},t))$, both depending on the same learned velocity $v_{\theta}$. Since the divergence term can be viewed as another output derived from the same velocity model, we can learn to distill both the flow trajectory and its corresponding divergence computation simultaneously within a single model. By jointly optimizing for both accurate few-step sampling and log-likelihood evaluation, we can potentially achieve a model that succeeds at both tasks.
@@
Based on these insights, we propose fast flow joint distillation (F2D2), a simple and modular framework for jointly learning fast sampling and fast log-likelihood evaluation in flow-based models. Our key idea is to leverage the flow map framework (Boffi et al., 2025a) to train a single model to predict both the sampling trajectory and cumulative divergence in parallel using a joint self-distillation objective, ensuring both outputs learn to skip the numerous steps in training. This makes F2D2 plug-and-play with any CNF-based few-step sampling method like shortcut models and MeanFlow, and requires only a new divergence prediction head alongside the existing velocity prediction. To our knowledge, F2D2 is the first method to enable accurate few-step log-likelihood evaluation in diffusion/CNF-based generative models, solving a long-standing limitation of these frameworks.
@@
We demonstrate that our method produces both calibrated likelihoods and high quality samples with few-step NFEs on image datasets CIFAR-10 (Krizhevsky et al., ) and ImageNet $64\times 64$ (Deng et al., 2009). We show that our F2D2 are compatible with and can be directly apply to pre-trained shortcut models, MeanFlow and a new distillation method we propose in this paper.
@@
As an application of our method, we introduce maximum likelihood self-guidance, a lightweight test-time intervention that uses rapid likelihood evaluation to optimize over generated samples, requiring only an additional forward and backward pass through the model. Remarkably, we show that F2D2 with maximum likelihood self-guidance instantiated with 2-step MeanFlow achieves lower FID than a 1024-step flow matching model of the same size on CIFAR-10. This proof of concept demonstrates the expanded algorithmic sandbox enabled by rapid likelihood evaluation.
@@ 2 Background
Flow Matching. Flow matching (Lipman et al., 2022; Albergo & Vanden-Eijnden, 2022; Liu et al., 2022; Albergo et al., 2023) is a scalable training method for generative modeling that learns a time-dependent velocity field $v_{\theta}:\mathbb{R}^{d}\times[0,1]\rightarrow\mathbb{R}^{d}$ to transport samples from a simple noise distribution $p_{0}$ to the data distribution $p_{1}$. Along a straight-line path $x_{t}=(1-t)x_{0}+tx_{1}$ that linearly interpolates between a noise sample $x_{0}\sim p_{0}$ and a data sample $x_{1}\sim p_{1}$, it models the evolution dynamic with the ordinary differential equation ODE:
@@
This objective encompasses diffusion models as a special case with different interpolation schemes (Gao et al., 2024). For sampling, it solves the ODE from $t=0$ to $t=1$ with numerical solvers like Euler or $\operatorname{dopri5}$, and typically requires 100-1000 NFEs for high quality samples.
Continuous Normalizing Flows and Likelihood Computation. A flow-matching model is a special case of a continuous normalizing flow (CNF) (Chen et al., 2018), which transports data from an initial distribution $x_{0}\sim p_{0}$ to an estimated distribution by integrating an ODE. In the case of flow-based models, this is precisely Eq.˜2.1. An advantage of the CNF formalism is the ability to explicitly compute likelihoods via the coupled system of ODEs:
@@
Likelihood evaluation is typically expensive, requiring both careful, finely-discretized integration of an ODE across time steps, and a computation of the divergence term whose exact computation (or the variance of its randomized estimator (Grathwohl et al., 2018)) scales at least linearly in ambient dimension $d$. Thus, likelihood evaluation is far more computationally burdensome than sampling.
Few-Step Flow-based Models. To address the computational expense of multiple ODE integrations in sampling, recent few-step flow-based models (Kim et al., 2023; Frans et al., 2024; Geng et al., 2025; Boffi et al., 2025a; b) learn to directly predict the outcome of integrating the ODE in Eq.˜2.1 using only a small number of function evaluations (NFEs). These methods can be viewed as sharing a common strategy of learning to predict the flow map of the underlying ODE.
@@ 3 Method
Our key insight is that we can apply few-step flow-based models to the ODE in Eq.˜2.3 which jointly parametrizes sampling and likelihood evaluation. Following Boffi et al. (2025a); Frans et al. (2024); Geng et al. (2025), we adopt a linear parametrization of the flow map:
@@
Proposition 3.1 (Flow Map Conditions (Boffi et al., 2025a)).
@@
Theoretical Justification. To justify the parameterization Eq.˜3.3, we recall a property denoted the tangent condition by Boffi et al. (2025a), leveraged by Kim et al. (2023); Geng et al. (2025) to recover the instantaneous velocity and divergence in the $s\to t$ limit:
@@ 3.2 Instantiating F2D2 with Shortcut and MeanFlow
Though our method is, in principle, compatible with any flow map-based method, we instantiate our formulation for Shortcut Models (Frans et al., 2024), based on the semigroup property, and for MeanFlow (Geng et al., 2025), based on the Eulerian equation.
@@ 3.2.1 Joint Shortcut: Shortcut-F2D2
Shortcut models (Frans et al., 2024) enforce the semigroup property (Proposition˜3.1 (c)) by applying it to the midpoint between timesteps $t$ and $s$. This amounts to the shortcut self-consistency loss:
@@ 3.2.2 Joint MeanFlow: MeanFlow-F2D2
Alternatively, MeanFlow (Geng et al., 2025) learns the time-averaged velocity $u_{\theta}(x,t,s)$ by enforcing the so-called MeanFlow identity $u_{\theta}(x_{t},s,t)=v(x_{t},t)+(s-t)\,\frac{d}{dt}u_{\theta}(x_{t},s,t)$, which we show in Corollary B.2 in Appendix B solves the Eulerian equation (Proposition˜3.1 (b)). To our knowledge, this is the first proof of this fact.
@@ 3.3 Practical Design Choices
Hutchinson Trace Estimator. Computing divergence terms $\operatorname{div}(v(x_{t},t))$ requires $O(d)$ backward passes for exact computation, making training prohibitively expensive for high-dimensional data. Following standard practice (Grathwohl et al., 2018; Lipman et al., 2022; Song et al., 2020), we employ the Hutchinson trace estimator $\operatorname{div}(v)\approx\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}[\epsilon^{\top}\nabla_{x}v\cdot\epsilon]$ which provides unbiased estimates with only $O(1)$ computational cost per training step.
Staged Training with Warm Start. Since divergence supervision depends on having accurate velocity predictions, we adopt a staged training approach. In practice, we pre-train the sampling velocity component $u_{\theta}$ alone using existing flow map distillation techniques, which provides a good initialization for joint training later. Optionally, we can also pre-train a teacher flow matching model $v_{\phi}$ that serves as a reliable source of divergence supervision, replacing the potentially noisy $\operatorname{div}(u_{\theta}(x_{t},t,t))$ with the more accurate $\operatorname{div}(v_{\phi}(x_{t},t))$ during joint distillation. When using a pre-trained teacher model or warm-starting, we replace the flow matching loss with a teacher matching loss, which will be described next.
@@
Adaptations to forward-only training. While shortcut models and MeanFlow are both consistent with the full flow-map formulation, their practical use (i.e. sampling) and the available pretrained checkpoints are forward-only: training and tuning assume $t<s$. Several implementation choices are exclusively designed for this regime (e.g., the specific parameters of MeanFlow’s logit-normal time sampling, shortcut models’ discretized time sampling schedule, and common logarithmical time-parameterizations used in EDM/CTM-style models). Extending training to explicitly cover both $t<s$ and $t>s$ would require re-deriving or re-tuning these design choices and can break the plug-and-play compatibility with existing pretrained models and algorithms.
In practice, we use a simple first-order approximation to reuse a forward-only model for backward likelihood integration. Let $\Delta x(x_{t},t,t+\Delta t)=\Delta t\,u_{\theta}(x_{t},t,t+\Delta t)$ denote the forward predicted displacement. We approximate the backward displacement by $\Delta x(x_{t},t,t-\Delta t)\approx-\Delta x(x_{t},t,t+\Delta t)$ and apply the same approximation to the log-density increment. This approximation is first-order accurate in $\Delta t$ and works well empirically at small-to-moderate step sizes. While there exist out-of-distribution inputs at $t=1$, in practice we find that the networks can generalize to these scenarios with forward-only training.
@@ 3.4 Application: Maximum Likelihood Self-Guidance with F2D2
Now that we have access to log-likelihood computation with few NFEs, we can explore various applications. One particularly interesting one is using the one-step divergence prediction (combined with the source distribution’s log-likelihood) as a pseudo-likelihood objective for inference-time optimization. Specifically, we can optimize the initial noise $x_{0}$ to improve sample quality before running the sampling procedure. This approach resembles reward-based initial noise optimization for one-step generation models (Eyring et al., 2024), except we do not require external reward models. Instead, we obtain the guidance signal from the model’s own likelihood prediction head – effectively performing self-guidance at inference time to improve sample quality. This maximum likelihood self-guidance sampling algorithm is described in Algorithm 2.